# Differentially Private Matchings

### v2

@@ 1.1.3 Node Privacy
[hsu2014private] show guarantees of the following form. Suppose each item has supply $s$, meaning that each item can be matched with $s$ bidders (let us call these $s$-matchings); if $s=\Omega(\log^{3.5}{n}/\varepsilon)$, then they give an algorithm which (implicitly) outputs a $(1+\eta)$-maximum matching while guaranteeing differential privacy. [^4] They also consider the weighted version of the problem and further generalizations where each bidder has gross substitutes valuations. Our techniques also extend to these cases with improved results, but we limit our discussion here to matchings as that is the focus of this work. They also show that $s=\Omega(1/\sqrt{\eta})$ is necessary to obtain any guarantees, leaving open the question of tightening these bounds. We substantially tighten the upper bound in Section 7, showing that $s=\Theta(\log{n}/\varepsilon)$ suffices:
@@
If the supply is at least $s\geq\Omega\left(\frac{\log(n)}{\eta^{4}\varepsilon}\right)$, Algorithm 5 outputs a $(1+\eta)$-approximate maximum (one-sided) $s$-matching with high probability.
1.1.4 Continual Release
Finally, all of our results translate to the continual release [DNPR10, CSS11] setting. In the graph continual release setting, edges are given as updates to the graph in a stream and the algorithm releases an output after each update in the stream. The continual release model requires the entire vector of outputs of the algorithm to be $\varepsilon$-differentially private. In this work, we focus on edge-insertion streams where updates contain edge insertions and edge-neighboring streams differ by one edge update.
We give results for two different types of streams in the continual release model. The first type of stream we consider is the arbitrary edge-order stream where the edges in the stream appear one by one in arbitrary order. In particular, in arbitrary edge-order streams, we show the following results for edge- and node-DP in Section 8. In our results below, we give bicriteria approximate solutions where an $(\psi,\phi)$-approximation is one with a $\psi$-multiplicative error and $\phi$-additive error. Our approximations hold, with high probability, for every update.
@@ 1.1.4 Continual Release
Let $\eta\in(0,1]$ and $\varepsilon\in(0,1)$. Algorithm 6 is an $\varepsilon$-edge DP algorithm in the arbitrary edge-order continual release model that outputs implicit $b$-matchings. Moreover, with high probability, (i) $b=O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)$and (ii) each implicit solution that has the size of a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum matching.
@@
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, and $\alpha$ be the arboricity of the input graph. There is an $\varepsilon$-node DP algorithm in the arbitrary edge-order continual release model that outputs implicit $b$-matchings. Moreover, with high probability, (i) $b=O\left(\frac{\alpha\log^{2}(n)}{\eta^{2}\varepsilon}+\frac{\log^{3}(n)}{\eta^{3}\varepsilon^{2}}\right)$and (ii) each implicit solution has the size of a
$\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta^{2}\varepsilon}\right)\right)$-approximate maximum matching.
The second type of stream we consider is the arbitrary adjacency-list order model [kallaugher2019complexity, goldreich2009proximity, czumaj2016relating, mcgregor2016better] in which nodes arrive in arbitrary order and once a node $v$ arrives as an update in the stream all edges adjacent to $v$ arrive in an arbitrary order as edge updates immediately after the arrival of $v$. In this model, we give the following results within Section 9.
@@
For $\varepsilon\in(0,1)$ and $b=O(\log(n)/\varepsilon)$, Algorithm 8 is an $\varepsilon$-edge DP algorithm in the arbitrary adjacency-list continual release model that outputs, with high probability, an implicit $b$-matching that has the size of a $(2,1)$-approximate maximum matching.
We summarize all of our results in Table 1.
@@
$\varepsilon$-LEDP ($n$ rounds) $O(\log(n)/\varepsilon)$ $2,(2+\eta)$ Upper Bound Theorem 1.3
@@
$\varepsilon$-LEDP ($O(\log n)$ rounds) $O(\log^{2}(n)/\varepsilon)$ $2$ Upper Bound Theorem 1.6
$\varepsilon$-node DP $O\left(\frac{\alpha\log(n)}{\eta\varepsilon}+{\frac{\log^{2}(n)}{\eta^{2}\varepsilon^{2}}}\right)$ $\left(2+\eta\right)$ Upper Bound Theorem 1.8
@@
Edge-Order $\varepsilon$-Edge DP $O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$ Upper Bound Theorem 1.11
@@
1.2 Technical Contributions of Independent Interest.
Our algorithms not only use ideas from parallel and distributed algorithms but also use several novel algorithmic tools, which may be of independent interest outside of the scope of this paper. In this section, we describe such novel contributions. In Section 1.3, we describe our contributions with respect to differentially private matchings.
Our lower bound proof relies on a symmetrization argument that enables us to prove the lower bound for a restricted class of highly symmetric algorithms and extend the hardness to all algorithms. To the best of our knowledge, this is the first application of symmetrization arguments towards DP lower bounds and may be of use outside of private matchings.
Our Public Vertex Subset Mechanism (Algorithm 4) allows us to privately release a set of candidate vertices per node. Each node can then intersect the released set with its private adjacency list to produce implicit vertex subset solutions. Our mechanism is an independent contribution that may be useful for other implicit graph problems beyond matching.
In order to implement our algorithms in the node-DP setting, we make novel use of matching sparsifiers for bounded-arboricity graphs [solomon2021local] by providing a “stable” implementation in the sense of [wagaman2024time] that allows us to apply our edge-DP algorithms in a black-box fashion on the sparsified input, while incurring a factor of the arboricity $\alpha$ in the approximation guarantees. While sparsification has been studied in the context of node privacy [wagaman2024time], previous utility guarantees typically depend on the maximum degree, which can be significantly bigger than the arboricity. [^5] The star graph on $n$ vertices has maximum degree $n-1$ but arboricity $\alpha=1$. We also prove a novel extension of the sparsifier to $b$-matchings, which may be of interest beyond private graph algorithms to the non-private graph algorithm setting. In general, bounded arboricity sparsifiers may be of use in node-DP beyond matchings.
1.3 Technical Overview
We now give an overview for some of the main ideas behind our results.
Explicit Solution Lower Bound. (Section 3)
At a very high level, our lower bound for explicit solutions (Theorem 1.1) uses an argument similar to the “packing arguments” of [DMN23] for minimum cut. However, instead of using a packing/averaging argument, we use a symmetry argument reminiscent of symmetry arguments in distributed computing (e.g. [kuhn2016local]) and semidefinite programming [dam2015semidefinite]. Namely, we argue that any DP algorithm for a collection of input graphs must be highly symmetric. Highly symmetric is defined in terms of the marginal probability of each non-existing edge; we show that the marginal probability of choosing any of these non-existent edges is identical. Thus, any DP algorithm that gives a solution with decent utility cannot satisfy differential privacy. And so, any differentially private algorithm must give a poor approximation on at least one of the graphs in our collection of input graphs. Our paper is, to the best of our knowledge, the first use of this type of symmetry argument in DP lower bounds.
1.3.1 Edge Privacy
LEDP algorithms. (Sections 4 and 5)
Our main algorithms (Theorem 1.3 and Theorem 1.6) take inspiration from greedy algorithm and distributed algorithms for maximal matching, although our proofs also hold for $b$-matchings. In particular, the traditional greedy algorithm for maximal matching iterates through the nodes in the graph in an arbitrary order and matches the nodes greedily whenever there exists an unmatched node in the current iteration. However, this algorithm is clearly not private because each matched edge reveals the existence of a true edge in the graph. Thus, we must have some way of releasing public subsets of vertices that include both true edges and non-edges.
Within this algorithm, we introduce the Public Vertex Subset Mechanism which may be of independent interest. In this mechanism, each node releases a public subset of nodes obtained via coin flips. Specifically, for each pair of nodes, we flip a set of coins with appropriate probability parameterized by subgraph indices $r$. For index $r$, each coin is flipped with probability $(1+\eta)^{-r}$. Then, a node $v$ publicly releases a subgraph index $r$ to determine a public subset of vertices consisting of coin flips which landed heads for each pair $\{v,u\}$ containing $v$. The value of $r$ determines the size of this public subset of vertices (in expectation). Then, each node $w$ takes the public subsets and intersects them with their private adjacency lists to determine a private subset of vertices only $w$ knows about. Thus, the public subset can be written on the billboard and the private subset is used to determine each node’s implicit answer. We believe this mechanism will be helpful for other graph problems that release implicit solutions.
To speed up our algorithm to run in polylogarithmic rounds (Theorem 1.6), we need to allow many vertices to simultaneously propose matches to their neighbors (rather than just one at a time). Intuitively, this can be done in ways similar to classical parallel and distributed algorithms for related problems like Boruvka’s parallel MST algorithm [Boruvka]. In these algorithms, nodes randomly choose whether to propose or to listen. Then, proposer nodes send proposals, receivers decide which of their proposals to accept, and receivers communicate the acceptances back to the proposers (see [DHIN19] for a recent example of this in distributed settings). However, the added LEDP constraint makes this process significantly more difficult. Note, for example, that even communicating “with neighbors” violates privacy, since the transcript would then reveal who the neighbors are! Hence, greater care must be taken regarding the messages each node transmits. Fortunately, by using techniques gleaned from our $n$-round algorithm, ideas from basic parallel algorithms, and improved analyses, our implementation goes through.
The crux of our $O(\log n)$ round distributed algorithm relies on our Public Vertex Subset Mechanism. We use this mechanism to release proposal sets on the part of the proposers and match sets on the part of the receivers. The proposers release public proposal sets that contain proposals to its neighbors to match. Using the Public Vertex Subset Mechanism, receivers can privately discern which nodes have proposed. Then, receivers also publish a public subset of nodes as their match sets. Using this mechanism, the proposers can then (privately) recognize which receivers accepted their proposals. Performing multiple rounds of this procedure leads to our desired Theorem 1.6. Our novel utility analyses depends on a novel noisy version of a charging argument, based on edge orientation, implicitly described in [khuller1994primal].
Implicit Solution Lower Bound.
Our lower bound on implicit solutions (Theorem 11.1) shows that we must violate the degree bounds in our returned matchings and so our bicriteria algorithms are reasonable. The proof is much more straightforward than the lower bound proof on explicit solutions. We start with a low-degree graph, and first use the DP guarantee over nearby graphs to argue that edges that do not exist must still be included with reasonably high probability (this is essentially a highly simplified version of the symmetry argument used in the previous lower bound). We then use group privacy to argue that there is a graph with logarithmic max degree (which is also logarithmically far from our starting graph) where many edges incident to the node of logarithmic degree must be included.
1.3.2 Node Privacy
General Graphs. (Section 6)
To achieve node privacy in the central model, we combine our edge privacy results with a “matching sparsifier” [solomon2021local]. Informally, a matching sparsifier of a graph $G$ is a subgraph $H$ such that 1) the maximum matching in $H$ is approximately as large as the maximum matching in $G$, and 2) the maximum degree of $H$ is at most $O(\alpha)$, where $\alpha$ is the arboricity of $G$. Note that the average degree in a graph with arboricity $\alpha$ is at most $O(\alpha)$, so such a sparsifier essentially turns the average degree into the maximum degree without harming the matching.
Given an algorithm to compute a matching sparsifier, it is obvious how to combine this with edge privacy to get node privacy: since the maximum degree in the sparsifier is at most $O(\alpha)$, basic group privacy implies that we can simply use our edge-private algorithm and incur an extra loss of at most an $O(\alpha)$ factor. In other words, if the maximum degree is small then node privacy is approximately the same as edge privacy, and a matching sparsifier guarantees small maximum degree.
So if we could compute and release a differentially private matching sparsifier, we would be done. Unfortunately, we show in Section 11.2 that this is impossible, by providing a strong lower bound against private matching sparsification (Theorem 11.2). However, we show that we can use these sparsifiers in a node-DP algorithm even though releasing them does not satisfy DP! Intuitively, this is because given two node-neighboring graphs, computing a sparsifier of them can bring the distributions of the outputs “closer together”. In particular, although these sparsifiers cannot be released publicly, they decrease the “edge-distance” between node-neighboring graphs, where the edge-distance is the number of edges that differ between neighboring graphs.
We extend our results to $b$-matchings using our novel arboricity sparsifiers for $b$-matchings in Section 10.2.
Improvement in Bipartite Graphs. (Section 7)
The starting point of our algorithm for node-private bipartite matchings is the same deferred acceptance type algorithm which [hsu2014private] base their private algorithm on. In their private implementation of the algorithm, they use continual counters [dwork2009complexity, CSS11] to keep track of the number of bidders each item is matched to. This is necessary for their implementation since their implicit solutions need to know the number of bidders an item is matched with at every iteration of the algorithm. We show that in a different private implementation of the algorithm, the counts of the number of bidders an item is matched to is only needed $O(1/\eta^{2})$ different times (where $\eta\in(0,1]$ is present in the $(1+\eta)$-approximation factor). Furthermore, a clever use of MAT to iteratively check if this count exceeds the threshold $s$ (where $s$ is the matching factor) suffices for the implementation, enabling us to substantially improve the error of the counts used.
1.3.3 Continual Release (Sections 8 and 9)
We extend all of our results to the continual release model under two different types of streams: edge-order streams and adjacency-list streams, both commonly seen in non-private streaming literature (see e.g. [kallaugher2019complexity, mcgregor2016better, mcgregor2005finding]). We consider edge-insertion streams where an update in the stream can either be $\bot$ (an empty update) or an insertion of an edge $\{u,v\}$. Edge-neighboring streams are two streams that differ in exactly one edge update. We also consider node-neighboring streams; node-neighboring streams are two streams that differ in all edges adjacent to any one node. We release a solution after every update over the course of $T=\operatorname{poly}(n)$ updates. [^6] Our algorithms extend to the case where $T=\omega(\operatorname{poly}(n))$ at the cost of factors of $\log(T)$. We focus on the case $T=\operatorname{poly}(n)$ as there are at most $O(n^{2})$ non-empty updates. The goal is to produce an accurate approximate solution for each update, with high probability.
We first adapt our LEDP and node-DP implicit $O(\log(n)/\varepsilon)$-matching algorithms to edge-order streams. These results are given in Theorem 1.11 and Theorem 1.12. The main idea behind our continual release algorithm is to use the sparse vector technique (SVT) to determine when to release a new solution at each timestep $t\in[T]$. Specifically, at timestep $t$, we check the current exact maximum matching size in the induced subgraph $G_{t}$ consisting of all updates up to and including $t$. If this matching size is greater than a $(1+\eta)$ factor (for some fixed $\eta\in(0,1]$) of our previously released solution, then we release a new solution. To release our new solution, we use our LEDP algorithm as a blackbox and pass $G_{t}$ into the algorithm. Then, we release the implicit solution that is the output of our LEDP algorithm. Since we can only increase our solution size $O(\log_{1+\eta}(n))$ times, we only accumulate an additional $O(\log(n)/\eta)$ factor in the error due to composition.
1.4 Related Work
Private Allocations.
[…]
Now, auction-based methods and the implicit solution concept of [hsu2014private] are inherently tied to bipartite graphs. While it is possible to reduce general matchings to a bipartite instance via random bipartitions [liu2024pointwise], this transformation incurs a factor of 2 in the multiplicative approximation factor and the obtained approximation ratio is only in expectation. This would imply a node-DP $(2+\eta)$-approximate $s$-matching in expectation for general graphs. Our novel solution concept allows us to handle general graphs, under edge-DP or node-DP assuming low arboricity, and obtain a $(2+\eta)$-approximate $s$-matching with high probability while reducing the lower bound to $s=\Omega(\log(n)/\varepsilon)$. Moreover, specializing our techniques to the bipartite case (with arbitrary arboricity) recovers the node-DP $(1+\eta)$-approximation ratio of [hsu2014private], with the same reduced bound on $s$.
We remark that a followup work [hsu2016jointly] studies a class of private convex programming problems that capture fractional allocations with supply $s$ under approximate differential privacy. Similar to our algorithm, they also output an $s^{\prime}$-matching with $s^{\prime}>s$. Unlike our results, they compute a fractional allocation and guarantee that the sum of constraint violations ($\sum_{j}s_{j}^{\prime}-s_{j}$) is bounded, whereas our algorithm returns an (integral) solution while ensuring each constraint violation is bounded ($\max_{j}s_{j}^{\prime}-s_{j}$). Thus, neither the privacy nor the utility guarantees are comparable to ours.
Private Matching Size.
Differentially private estimations of the size of the maximum matching have been recently studied in various privacy models including the continual release model [FHO21, wagaman2024time, raskhodnikova2024fully, epasto2024sublinear] and the sublinear model [BGM22]. In the standard central DP model, computing an approximate solution for the size of the maximum matching in the input graph with $O(\log(n)/\varepsilon)$ additive error is trivial since the sensitivity of the size of the maximum matching is $1$ and so we can use the Laplace mechanism to obtain such an estimate.
Implicit Solutions.
Implicit solutions in the graph setting were first introduced by [GuptaLMRT10] for various problems such as set cover and vertex cover. An implicit solution is one where the released information from a differentially private algorithm is $\varepsilon$-DP or $\varepsilon$-LEDP but the released solution does not immediately solve the problem. In particular, the released solution is an implicit solution from which each individual node can compute the explicit solution for itself. In the case of our matching algorithms, the explicit solution computed by every node is the set of edges in the matching that are adjacent to it.

### v3

@@ 2.4 Utility Improvements in Bipartite Graphs over [hsu2014private] (Section 8)
[hsu2014private] show guarantees of the following form. Suppose each item has supply $s$, meaning that each item can be matched with $s$ bidders (let us call these $s$-matchings); if $s=\Omega\left(\frac{\log^{3.5}n}{\varepsilon}\right)$, then they give an algorithm which (implicitly) outputs a $(1+\eta)$-maximum matching while guaranteeing differential privacy. [^5] They also consider the weighted version of the problem and further generalizations where each bidder has gross substitutes valuations. Our techniques also extend to these cases with improved results, but we limit our discussion here to matchings as that is the focus of this work. They also show that $s=\Omega(\nicefrac{{1}}{{\sqrt{\eta}}})$ is necessary to obtain any guarantees, leaving open the question of tightening these bounds. We substantially tighten the upper bound in Section 8, showing that $s=\Theta\left(\frac{\log{n}}{\varepsilon}\right)$ suffices:
@@
If the supply is at least $s\geq\Omega\left(\frac{\log(n)}{\eta^{4}\varepsilon}\right)$, Algorithm 5 is $\varepsilon$-node DP and outputs a $(1+\eta)$-approximate maximum (one-sided) $s$-matching with high probability.
The starting point of our algorithm for node-private bipartite matchings is the same deferred acceptance type algorithm which [hsu2014private] base their private algorithm on. In their private implementation of the algorithm, they use continual counters [dwork2009complexity, CSS11] to keep track of the number of bidders each item is matched to. This is necessary for their implementation since their implicit solutions need to know the number of bidders an item is matched with at every iteration of the algorithm. We show that in a different private implementation of the algorithm, the counts of the number of bidders an item is matched to is only needed $O(1/\eta^{2})$ different times (where $\eta\in(0,1]$ is present in the $(1+\eta)$-approximation factor). Furthermore, a clever use of the Multidimensional AboveThreshold mechanism (Algorithm 1) to iteratively check if this count exceeds the threshold $s$ (where $s$ is the matching factor) suffices for the implementation, enabling us to substantially improve the error of the counts used.
2.5 Continual Release (Section 9)
We extend all of our results to the continual release model [DNPR10, CSS11] under two different types of streams: edge-order streams and adjacency-list streams, both commonly seen in non-private streaming literature (see e.g. [kallaugher2019complexity, mcgregor2016better, mcgregor2005finding]). In the graph continual release setting, edges are given as updates to the graph in a stream and the algorithm releases an output after each update in the stream. The continual release model requires the entire vector of outputs of the algorithm to be $\varepsilon$-differentially private. We consider edge-insertion streams where an update in the stream can either be $\bot$ (an empty update) or an insertion of an edge $\{u,v\}$. Edge-neighboring streams are two streams that differ in exactly one edge update. We also consider node-neighboring streams; node-neighboring streams are two streams that differ in all edges adjacent to any one node. We release a solution after every update over the course of $T=\operatorname{poly}(n)$ updates. [^6] Our algorithms extend to the case where $T=\omega(\operatorname{poly}(n))$ at the cost of factors of $\log(T)$. We focus on the case $T=\operatorname{poly}(n)$ as there are at most $O(n^{2})$ non-empty updates. The goal is to produce an accurate approximate solution for each update, with high probability.
We first adapt our LEDP and node-DP implicit $O(\log(n)/\varepsilon)$-matching algorithms to edge-order streams. These results are given in Corollary 9.1 and Theorem 2.10. The main idea behind our continual release algorithm is to use the sparse vector technique (SVT) to determine when to release a new solution at each timestep $t\in[T]$. Specifically, at timestep $t$, we check the current exact maximum matching size in the induced subgraph $G_{t}$ consisting of all updates up to and including $t$. If this matching size is greater than a $(1+\eta)$ factor (for some fixed $\eta\in(0,1]$) of our previously released solution, then we release a new solution. To release our new solution, we use our LEDP algorithm as a blackbox and pass $G_{t}$ into the algorithm. Then, we release the implicit solution that is the output of our LEDP algorithm. Since we can only increase our solution size $O(\log_{1+\eta}(n))$ times, we only accumulate an additional $O(\log(n)/\eta)$ factor in the error due to composition.
Our results are given formally below. We give bicriteria approximate solutions where an $(\psi,\phi)$-approximation is one with a $\psi$-multiplicative error and $\phi$-additive error. Our approximations hold, with high probability, for every update.
@@ 2.5 Continual Release (Section 9)
For $\varrho,\eta\in(0,1)$ and $b\geq\frac{(1+\eta)^{2}}{1-\eta}b^{\prime}+O\left(\frac{\log^{2}(n)}{\varrho\eta^{2}\varepsilon}\right),$ there is an $\varepsilon$-edge DP algorithm for the arbitrary edge-order continual release model that outputs a sequence of implicit $b$-matchings in the billboard model such that, with high probability, each $b$-matching is a $\left(2+\varrho,\left(\frac{\log^{2}(n)}{\varrho\varepsilon}\right)\right)$-approximation with respect to the maximum $b^{\prime}$-matching size.
Similar to our static edge-DP algorithms, there are two regimes of particular interest.
  • If $b=\Omega\left(\frac{\log^{2}(n)}{\varrho\varepsilon}\right)$, there is an $\varepsilon$-edge DP algorithm that outputs a sequence of implicit $b$-matchings with the size of a $\left(2+\varrho,\left(\frac{\log^{2}(n)}{\varrho\varepsilon}\right)\right)$-approximate 1-matching (Corollary 9.1).
  • If $b=\Omega\left(\frac{\log^{2}(n)}{\eta^{4}\varepsilon}\right)$, there is an $\varepsilon$-edge DP algorithm that outputs a sequence of implicit $b$-matchings that are $\left(2+\varrho,\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate $b$-matchings (Theorem 9.2). Note that we circumvent the need for a bicriteria approximation.
Our continual release algorithm can also be implemented under node-DP by implementing the appropriate arboricity sparsifier in the online setting.
@@
Let $\eta\in(0,1]$, $\varepsilon\in(0,1)$, $\alpha$ be the arboricity of the input graph, and
$b=\Omega\left(\frac{\alpha\log^{2}(n)}{\eta^{2}\varepsilon}+\frac{\log^{3}(n)}{\eta^{2}\varepsilon^{2}}+\frac{b^{\prime}\log^{2}(n)}{\eta\varepsilon}\right)\,.$
There is an $\varepsilon$-node DP algorithm for the arbitrary edge-order continual release model that outputs a sequence of implicit $b$-matchings in the billboard model such that, with high probability, each $b$-matching has the size of a $\left(2+\eta,O\left(\frac{b^{\prime}\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate maximum $b^{\prime}$-matching.
As a special case, we have the following corollary:
  • If $b=\Omega\left(\frac{\alpha\log^{2}(n)}{\eta^{2}\varepsilon}+\frac{\log^{3}(n)}{\eta^{2}\varepsilon^{2}}\right)$, there is a $\varepsilon$-node DP algorithm that outputs a sequence of $b$-matchings that are the size of a $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$-approximate 1-matching (Corollary 9.3).
The second type of stream we consider is the arbitrary adjacency-list order model [kallaugher2019complexity, goldreich2009proximity, czumaj2016relating, mcgregor2016better] in which nodes arrive in arbitrary order and once a node $v$ arrives as an update in the stream all edges adjacent to $v$ arrive in an arbitrary order as edge updates immediately after the arrival of $v$. In this model, we give the following results within Section 9.3.
@@
For $\varepsilon\in(0,1)$ and $b=O(\log(n)/\varepsilon)$, Algorithm 8 is an $\varepsilon$-edge DP algorithm in the arbitrary adjacency-list continual release model that outputs a sequence of implicit $b$-matchings such that, with high probability, each $b$-matching has the same guarantee as Theorem 2.2 with the addition of an additive error of $1$ per update.
Our quantitative results are summarized in Table 1. For simplicity, we state all approximation guarantees and lower bounds with respect to the maximum matchings, whereas our algorithms can also produce approximations with respect to the optimal $b$-matching.
@@
$\varepsilon$-LEDP ($n$ rounds) $O(\log(n)/\varepsilon)$ $2,(2+\eta)$ Upper Bound Corollary 5.1
@@
$\varepsilon$-LEDP ($O(\log n)$ rounds) $O(\log^{2}(n)/\varepsilon)$ 22 Upper Bound Theorem 2.3
$\varepsilon$-node DP $O\left(\frac{\alpha\log(n)}{\eta\varepsilon}+{\frac{\log^{2}(n)}{\eta^{2}\varepsilon^{2}}}\right)$ $\left(2+\eta\right)$ Upper Bound Corollary 7.1
@@
Edge-Order $\varepsilon$-Edge DP $O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)$ $\left(2+\eta,O\left(\frac{\log^{2}(n)}{\eta\varepsilon}\right)\right)$ Upper Bound Corollary 9.1